# Almost commuting self-adjoint operators and measurements

### v3

@@ 4 Multiple self-adjoint operators
Fix For any there exists satisfying the following: Suppose that is a unital -algebra, is a separable ideal of real rank zero, is the quotient map such that is purely infinite simple, and with () such that the -tuple has a -near-spectrum and has -spectrum Then there are mutually orthogonal non-zero projections a -dense subset in and a c.p.c. map where such that
| (e 4.128) | |||
| (e 4.129) |
@@
Let Choose Since is a -near-spectrum of there is a unital c.p.c. map such that
| (e 4.130) |
Since is a -spectrum of there exists, by Corollary 3.7, a compact subset such that and a unital monomorphism such that
| (e 4.131) |
@@
| (e 4.148) | |||||
| (e 4.149) |
Hence, using the inequality above and (e 4.130),
| (e 4.150) |
@@
Let be a unital -algebra. Let and be maps, and We say satisfies condition if has --cancellation with for all -stable rank -divisible rank for all and for all
@@
We will apply the following theorem which is a variation of that of Theorem 3.1 of [10].
@@
Let where is a compact subset. Let be maps, and
For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebra satisfying condition such that for any compact metric space and are unital c.p.c. maps such that
| (e 4.151) |
then there are a unitary and a homomorphism with finite dimensional image such that
| (e 4.152) |
Proof.
Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently multiplicative c.p.c. maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
| (e 4.153) |
| (e 4.154) |
Theorem 4.6.
Let and Let and be maps and There exists satisfying the following: Suppose that is a unital -algebra such that for any compact metric space () satisfying condition and is an essential ideal which has real rank zero such that is purely infinite simple, for any non-zero projection and suppose that with () such that
@@
where and where is the quotient map. Then there are such that
@@
Fix Let Fix a finite -dense subset such as for some large (as in 2.6). Assume that there are many points in There are at most many compact subsets of such that, for any compact subset there exists such that
@@
We will apply Theorem 4.5. Let (in place of ) be given by Theorem 4.5 for and (in place of ), Define
@@
| (e 4.159) |
By the choice of (and ) and, applying Proposition 2.11, and are non-empty set (see also (e 2.15)) and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ), we have (by applying Proposition 2.15),
| (e 4.160) |
By the choice of applying Lemma 4.2, we obtain mutually orthogonal nonzero projections a -dense subset a c.p.c. map such that (with )
| (e 4.161) | |||
| (e 4.162) |
and Put for all
@@
By the choice of and (e 4.162), as and applying Theorem 4.5, we obtain a unital homomorphism with finite dimensional range and a unitary such that
| (e 4.163) |
@@
| (e 4.164) |
| (e 4.165) |
where
By the assumption that and (e 4.160), is -dense in Note that
| (e 4.166) |
Hence is -dense in Let
| (e 4.167) | |||||
| (e 4.168) |
By induction, since is -dense in we obtain such that and where
Put Define
| (e 4.169) |
Then
| (e 4.170) |
In we have a partial isometry such that
Put Then
| (e 4.171) |
It follows that
| (e 4.172) | |||||
| (e 4.173) |
@@ 5 A pair of almost commuting self-adjoint operators on Hilbert modules
It follows from Lemma 4 of [23] that, for any projection has (IR).
@@
| (e 5.180) |
whenever
(Note that the identity in (e 5.180) also implies that )
@@
| (e 5.182) |
Put and Then is a normal element in and Suppose that such that Then
| (e 5.183) |
Hence
| (e 5.184) |
@@
Let There exists satisfying the following: Suppose that is a unital -algebra with an essential ideal which is a -unital -algebra with real rank zero and with property (IR) such that is purely infinite simple. Suppose that with () such that
| (e 5.185) |
for all where is the quotient map. Then, if either has real rank zero, or is path connected, there are such that
| (e 5.186) |
@@
Choose Let be given by Proposition 2.15 for (in place of ) and Choose Let be given by Proposition 2.16 for (in place of ). Let be given by Theorem 5.2 for (in place of ).
@@
Now assume that satisfy the assumption for Put
@@
| (e 5.187) |
for those for which where and In particular, has an On the other hand, by the choice of we have
| (e 5.188) |
@@
| (e 5.189) |
To see the claim, we note that, by (e 5.188), Then, by (e 5.187) and by the assumption (e 5.185),
| (e 5.190) |
@@
| (e 5.194) |
Hence, by the second part of (e 5.192),
| (e 5.195) |
If and (i.e., in the case (ii)), by the second part of (e 5.192) and the claim above, one computes that
| (e 5.196) |
@@
| (e 5.203) |
Moreover, by the second part of (e 5.202) and (e 5.201), and then, by the first part of (e 5.202), (e 5.201), (e 5.203), we obtain
| (e 5.204) | |||
| (e 5.205) | |||
| (e 5.206) |
@@
| (e 5.207) |
It follows from (e 5.205) and (e 5.206), and by the assumption on and that
| (e 5.208) | |||
| (e 5.209) |
@@
For (iii), let and As in 4.7 (the proof of Theorem 1.2), by Kasparov’s theorems, and is a -unital hereditary -subalgebra of a -unital simple -algebra of real rank zero and stable rank one. Hence is also a simple -algebra of real rank zero and stable rank one, whence has property (IR). If is unital, then Therefore the conclusion follows from Theorem 4.4 of [8]. Otherwise, by the assumption, has continuous scale (see of [27]). Hence, by Corollary 3.3 of [27], is a simple purely infinite simple -algebra. Thus Theorem 5.3 applies. ∎
@@
The next proposition provides a short way to see the second condition in (e 5.185) is necessary. We also provide a direct proof.
@@
| (e 5.217) |
Note that (e 5.217) also implies that is invertible.
@@
| (e 5.221) |
Now if then, by (e 5.220), Since is normal,
| (e 5.222) |
Hence, by (e 5.221),
| (e 5.223) |
@@
There is a considerable overlap between the results in this section and that of Theorem 1.1 of [17] even though they have different proofs. However, Theorem 1.1 of [17] provides a quantitative estimate of the distance from to the set of normal elements with finite spectrum (perhaps the quantity is not ready to be computed) while Theorem 5.3 is a conceptional result. In the special case of (i) of Theorem 1.3, (defined in [17]) is equivalent to the condition that for all
On the other hand, computationally speaking, -synthetic-spectrum could be tested. Note that a -synthetic-spectrum consists of finitely many balls and there are only finitely many bounded connected components in its complement. This allows one to avoid the situation of infinitely many “holes” in the spectrum. Note that the index in each of these “holes” remains constant. So one only needs to test one point in each of these “holes”. One also notes that the second condition in (e 5.185) does not imply
It should also be noted that, when or a normal element ( is Theorem 5.3) may not be approximated by normal elements with finite spectrum even under the assumption that for all
In Theorem 1.1 of [17], seems depending on a given operator The constant in Theorem 5.3 does not depend on nor even -algebras (as long as satisfies the conditions in Theorem 5.3). Finally, let us note that when a unital -algebra has (IR), then, by Theorem 4.4 of [8], a pair of almost commuting self-adjoint elements and is close to a pair of commuting ones even though may be large (see also the example below).
@@
Note that is purely infinite, has property (IR). Since every projection in lifts to a projection in (see Lemma 4.1). It follows from Corollary 3.16 of [3] that has real rank zero. Hence Theorem 5.3 applies in this case. However, from the six-term exact sequence in -theory, one computes that Denote by the set of normal elements of and the set of those elements in which has finite spectrum, respectively. Then the closure Let be a non-zero projection. Since is either purely infinite or stable rank one, Thus there is a unitary with Let Then is a normal element with Hence for any Since and one sees easily that In other words, Theorem 1.1 in [17] could not be applied to determine when and are almost commuting (see also Theorem in [23]).
@@ 6 Epilogue
In this last section, we will address the relation between the condition of vanishing Fredholm index and that of closeness of the (synthetic)-spectrum and essential (synthetic)-spectrum (see Corollary 6.2 and 6.3). We will also give an example (Proposition 6.4) which shows the complexity behind the failure of a higher dimensional version of Theorem 1.3. A final remark on the Mumford original question related to quantum theory and measurements, and their relation with the second condition in Theorem 1.1 is also included.
@@
The next corollary shows that, when the second condition in (e 4.156) is stronger than that of vanishing index in Theorem 1.3.
@@
We choose such that the conclusion of Proposition 2.15 holds for Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (e 6.226) holds. Put Thus (applying Proposition 2.11) and are all non-empty sets. By Proposition 2.15,
| (e 6.228) |
@@
| (e 6.232) |
So, if By [5], in case (i) and (ii), since is stable, For case (iii), follows from Lemma 3.3 of [19] (note that, proved in [18], if has real rank zero, then ). Therefore
| (e 6.233) |
@@
| (e 6.234) |
In general, we have the following statement. This follows immediately from Theorem 4.6. With the same spirit as that of the proof of Corollary 6.2, one could provide a direct proof without mentioning Theorem 4.6. However, since the point has been made in Corollary 6.2, we refrain ourselves from placing further -theory burden on the reader who might not be familiar with -theory.
@@
We now present an example which shows that an -dimensional version of part (i) of Theorem 1.3 is false (). Note that, in the next proposition, the triple has a single point in the essential -spectrum. Hence the -theory obstruction vanishes in
@@
Choose Identify as a finite rank () projection. Put Define by
| (e 6.243) |
Choose
| (e 6.244) |
@@
| (e 6.257) |
This contradicts (e 6.239). Therefore there is no homomorphisms such that (e 6.249). In other words, there is such that
| (e 6.258) |
where the infimum is taken among all homomorphisms Combining this with (e 6.246), the proposition follows. ∎
@@
This remark was added to the original version of the paper after the author fortunately received a copy of Chapter 14 of a very exciting new book of David Mumford [34] graciously sent by the author. Some terminologies are taken from there.
@@
In quantum theory, one may view macroscopic observables as some self-adjoint operators on a Hilbert space which we assume to be infinite dimensional and separable. Let be the unital -subalgebra generated by and (let us assume for convenience here). Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonally) a finite dimensional subspace vanish. In other words, outsiders would see nothing. Let us assume that none of these observables collapse this way (see also Remark 6.5). In other words, we assume that none of ’s are compact.
@@
Given a state which is defined by a unit vector the expected value when measuring an observable given by an operator is (see p.178 of [34]). “Approximately Macroscopically Unique” states introduced by Mumford (see Chapter 14 of [34]), or AMU states, are “near eigenvectors”. Set, for a small
@@
where As pointed out in [34], for AMU to be non-empty, it is natural to assume that the commutators have small bound. If the question in the introduction (see (e 1.1) and (e 1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one. Consequently, AMU does exist. Observed by Mumford, the condition that the commutators are small is not quite enough (Proposition 6.4 further shows the complexity of the problem) in general.
@@
In an infinite dimensional separable Hilbert space a self-adjoint operator such as may not have any eigenvalues (but have approximate eigenvalues). However, if one allows non-normal states of or that of then does have eigenstates and eigenvalues in the atomic representation of or (see, for example, Proposition 4.3.10 of [35]). In particular, one should consider those states of which vanish on the ideal of compact operators. If a point in the unit cube of is expected to be a “joint eigenvalue”, then one may use the test function defined in 2.7 to test it. In order to have (for small ), one only needs one vector state (with a unit vector ) such that
| (e 6.259) |
@@
One may expect this can be repeatedly measured. Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (e 6.259) also holds for (in place of ), and this persists. Then we obtain a sequence of () such that (e 6.259) holds for each (in place of ). Note that unit vector states given by has limit points which give states of vanishing on and Note that there are still unknown (approximate) eigenvalues outside any finite dimensional subspaces of
@@
Let be the quotient map and be a state of the corona algebra (or that of ). Then the quantum state might better reflect this situation. Indeed, one may expect that, at least for one such quantum state one still has
| (e 6.260) |
@@
Measurements given by states of the form may be called outside measurements. By the Gelfand-Naimark construction, this quantum state is also given by a vector state (the Hilbert space from the atomic representation of or ). In other words,
| (e 6.261) |
@@
If whenever (e 6.259) holds, inequality (e 6.260) holds for some quantum states then the second condition in (e 1.3) always satisfies (in fact, in this case ).
@@
Let us, for a moment, call a unit vector state for which (e 6.259) holds a synthetic-AMU state and a synthetic-joint-expected value. The state for which (e 6.260) holds may be called an essential synthetic-AMU and (for which (e 6.260) holds) an essential synthetic-joint-expected value. Note that it is easy to see an AMU state (with sufficiently small ) is an ()-synthetic-AMU state. Moreover, by Proposition 2.11, the set of synthetic-AMU states is non-empty, whenever is small (independent of ). The second condition in Theorem 1.1 then may be stated as follows:
@@
(C): all synthetic-joint-expected values are in a sufficiently small neighborhood of the set of essential synthetic-joint-expected values.
@@
This seems to suggest that we could put a condition on the macroscopic system: any local measurement should not be too far off from an outside (or a long-term measurements along an orthonormal basis of the Hilbert space) measurement. Under this condition, when commutators tend to zero, by Theorem 1.1, the quantum system may be approximated by a classical system where observables commute.

### v4

@@ 4 Multiple self-adjoint operators
Fix For any there exists satisfying the following: Suppose that is a unital -algebra with an ideal such that:
1. is separable and has real rank zero;
2. the quotient is purely infinite ans simple.
3. with ();
4. the -tuple has a -near-spectrum and has a -spectrum Then, there are mutually orthogonal non-zero projections a -dense subset in and a c.p.c. map where such that
| (e 4.133) | |||
| (e 4.134) |
@@
Let Choose Since is a -near-spectrum of by Definition 2.9, there is a unital c.p.c. map such that
| (e 4.135) |
Since is a -spectrum of there exists, by Definition 2.9 and by Corollary 3.7, a compact subset such that and a unital monomorphism such that
| (e 4.136) |
@@
| (e 4.154) | |||||
Hence, using the inequality above and (e 4.135),
| (e 4.155) |
@@
Let be a unital -algebra. Let and be maps, and We say satisfies condition if has --cancellation, -stable rank -divisible rank for all and for all
@@
We will apply the following theorem which is a variation of Theorem 3.1 of [10].
@@
Fix Let be maps, and
For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebra such that:
1. satisfies condition ;
2. for any compact metric space
Let be a compact subset of and let be unital c.p.c. maps such that
@@
Then, there are a unitary (for some ) and a homomorphism with finite dimensional image, such that
@@
Let us first fix a compact subset Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently multiplicative c.p.c. maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
@@
| (e 4.159) |
However, since generates as -algebra, one obtains such that when (e 4.156) holds, then
| (e 4.160) |
So Theorem 3.1 of [10] applies and the theorem would follow if we fix first.
To find a common and indepedent of we note that there are finitely many compact subsets of such that, for any non-empty compact subset there is such that
| (e 4.161) |
(see 2.14). Let Then, for any non-empty compact subset of there is such that
| (e 4.162) |
When is given, by the first part of the proof, we obtain and such that the conclusion of the first part holds for Choose and
Now suppose and are unital c.p.c. maps satisfying the assumption (e 4.156). Suppose that
| (e 4.163) |
Define by (this makes sense as ), Therefore
| (e 4.164) |
and Hence, with the choice of we obtain homomorphism with finite dimensional range and a unitary such that, for
| (e 4.165) |
We may write where () and is a set of mutually orthogonal projections in Since for each we may choose such that
| (e 4.166) |
Define by
| (e 4.167) |
Note that
| (e 4.168) |
Hence, by (e 4.165), for
| (e 4.169) |
Theorem 4.6.
Let and Let and be maps and There exists satisfying the following:
Suppose that is a unital -algebra such that:
1. satisfies condition
2. for any compact metric space ();
3. is an essential ideal which has real rank zero such that is purely infinite simple;
4. for any and any non-zero projection we have
Suppose further that with () satisfy :
| (e 4.170) |
(recall Definition 2.14), where and where is the quotient map.
Then there are such that
| (e 4.171) |
Proof.
@@
| (e 4.172) |
By the choice of (and ) and, applying Proposition 2.11, we obtain non-empty sets and (see also (e 2.15)), and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ), since we have (by applying Proposition 2.15),
| (e 4.173) |
By the choice of (and ), applying Lemma 4.2, we obtain mutually orthogonal nonzero projections a -dense subset a c.p.c. map such that (with )
| (e 4.174) | |||
| (e 4.175) |
and Put
@@
By the choice of and (e 4.175), as and applying Theorem 4.5, we obtain a unital homomorphism (for some integer ) with finite dimensional range and a unitary such that
| (e 4.176) |
| (e 4.177) |
where Put Note that, without loss of generality, by replacing by and adding a projection to each if neseccarily, we may assume that
Define by
| (e 4.178) |
Let be a chosen matrix unit for In what follows, we identify with and above with where In particular, we view
Recall that is -dense in By the assumption that and (e 4.173), we conclude that is -dense in Note that
| (e 4.179) |
Hence is -dense in Let
| (e 4.180) | |||||
| (e 4.181) |
By induction, since is -dense in we obtain mutually disjoint subsets such that and where
Put Define
| (e 4.182) |
Then
| (e 4.183) |
Recall that, for any nonzero projection Thus, in viewing we have a partial isometry such that
Then put Then
| (e 4.184) |
It follows that
| (e 4.185) | |||||
| (e 4.186) |
@@
| (e 4.187) |
for all Then, by (e 4.176), (e 4.178), (e 4.183), (e 4.186) and (e 4.174), for
| (e 4.188) |
@@ 5 A pair of almost commuting self-adjoint operators on Hilbert modules
It follows from Lemma 4 of [23] that, for nonzero any projection has (IR).
@@
| (e 5.193) |
(recall Definition 2.4), whenever
(Note that the identity in (e 5.193) also implies that )
@@
| (e 5.195) |
Put and Then is a normal element in and Moreover
| (e 5.196) |
Suppose that such that Then
| (e 5.197) |
Hence (also by (e 5.196))
| (e 5.198) |
@@
Let There exists satisfying the following: Suppose that is a unital -algebra with an essential ideal such that:
1. is -unital, has real rank zero and satisfies property (IR);
2. the quotient is purely infinite simple.
Let with () such that:
| (e 5.199) |
for all where is the quotient map.
Then, if either:
has real rank zero, or
is path connected,
there are such that
| (e 5.200) |
@@
Choose Let be given by Proposition 2.15 for (in place of ) and Choose
Let be given by Theorem 5.2 for (in place of ). Put
Let be given by Proposition 2.16 for (in place of ).
@@
Now assume that satisfy the assumption for By Proposition 2.16, Put (see (e 2.15))
@@
| (e 5.201) |
for those for which where and In particular, the pair has an -spectrum We may view it as an On the other hand, by the choice of and applying Proposition 2.15, we have
| (e 5.202) |
@@
| (e 5.203) |
To see the claim, we note that, by (e 5.202), Then, by (e 5.201) and by the assumption (e 5.199),
| (e 5.204) |
@@
| (e 5.208) |
Hence, by the second part of (e 5.206),
| (e 5.209) |
If and (i.e., in the case (ii)), by the second part of (e 5.206) and the claim above, one computes that
| (e 5.210) |
@@
| (e 5.217) |
Moreover, by the second part of (e 5.216) and (e 5.215), and then, by the first part of (e 5.216), (e 5.215), (e 5.217), we obtain
| (e 5.218) | |||
| (e 5.219) | |||
| (e 5.220) |
@@
| (e 5.221) |
It follows from (e 5.219) and (e 5.220), and by the assumption on and that
| (e 5.222) | |||
| (e 5.223) |
@@
For (iii), let and As in 4.7 (the proof of Theorem 1.2), by Kasparov’s theorems, and is a -unital hereditary -subalgebra of a -unital simple -algebra of real rank zero and stable rank one. Hence is also a simple -algebra of real rank zero and stable rank one, whence has property (IR). If is unital, then Therefore the conclusion follows from Theorem 4.4 of [8]. Otherwise, by the assumption, is simple. It follows that has continuous scale (see of [27]). Hence, by Corollary 3.3 of [27], is a simple purely infinite simple -algebra. Thus Theorem 5.3 applies. ∎
@@
The next proposition provides a short way to see the second condition in (e 5.199) is necessary. We also provide a direct proof.
@@
| (e 5.231) |
Note that (e 5.231) also implies that is invertible.
@@
| (e 5.235) |
Now if then, by (e 5.234), Since is normal,
| (e 5.236) |
Hence, by (e 5.235),
| (e 5.237) |
@@
Let us compare the results in this section and Theorem 1.1 of [17]:
While there is significant overlap between the results in this section and Theorem 1.1 of [17], the proofs are quite different. Theorem 1.1 of [17] provides a quantitative estimate for the distance from to the set of normal elements with finite spectrum (though the explicit computation of may not be straightforward). In contrast, Theorem 5.3 here is a more conceptual result. In the special case of Theorem 1.3 (i), the condition (defined in [17]) is equivalent to for all
Key Differences and Limitations:
However, when or a normal element (as in Theorem 5.3) may not be approximable by normal elements with finite spectrum—even if for all In such cases, (from Theorem 1.1 of [17]) can be large. For instance, if is a unitary with in then rendering Theorem 1.1 of [17] inapplicable (see Example 5.7 for details).
Moreover, if a unital -algebra has property (IR), then by Theorem 4.4 of [8], any pair of almost commuting self-adjoint elements and is close to a commuting pair – regardless of the size of (see also Example 5.7 below).
Computational Aspects of -Synthetic Spectrum:
From a computational standpoint, the -synthetic spectrum is tractable: it consists of finitely many balls, and its complement has only finitely many bounded connected components. Since the index remains constant in each such component, it suffices to test one point per bounded components. This avoids the complication of infinitely many “holes” in the spectrum
Further points:
The second condition in (e 5.199) does not imply
@@
Let be the quotient map. Note that is purely infinite, has property (IR). Since every projection in lifts to a projection in (see Lemma 4.1). It follows from Corollary 3.16 of [3] that has real rank zero. Hence Theorem 5.3 applies in this case. However, from the six-term exact sequence in -theory, one computes that Denote by the set of normal elements of and the set of those elements in which has finite spectrum, respectively. Then the closure
Let be a non-zero projection. Since is either purely infinite or stable rank one, Thus there is a unitary with Let Then is a normal element with Hence for any (so Theorem 1.3 (iii) applies).
Then Choose Put Suppose that with Then, for any
| (e 5.238) |
It follows that In other words,
Let be the disk in with center at and radius Then Choose such that if if if Define such that (). Note that and Then, there exists such that, for any pair of normal elements with and with then
| (e 5.239) |
Put
Claim: but
In fact, suppose otherwise there is such that
| (e 5.240) |
Then, since and
| (e 5.241) |
Hence and Therefore is a normal partial isometry. Moreover,
| (e 5.242) |
There exists a unitray such that
| (e 5.243) |
Also
Note that is a unitary in with finite spectrum, hence This contradict the fact that in which proves the claim.
@@ 6 Epilogue
In this final section, we examine the connection between the vanishing Fredholm index condition and the closeness of the (synthetic) spectrum to the essential (synthetic) spectrum (Corollaries 6.2 and 6.3). We also present an example (Proposition 6.4) where a higher-dimensional analogue of Theorem 1.3 fails– highlighting the obstructions and subtleties behind this failure. Finally, we conclude with a remark on Mumford’s original question regarding quantum theory and measurements, relating it to the second condition in Theorem 1.1.
@@
The next corollary shows that, when the second condition in (e 4.170) is stronger than that of vanishing index in Theorem 1.3.
@@
We choose such that the conclusion of Proposition 2.15 holds for (in place of ). Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (e 6.246) holds. Put Thus (applying Proposition 2.11) and are all non-empty sets. By Proposition 2.15,
| (e 6.248) |
@@
| (e 6.252) |
So, if then By [5], in case (i) and (ii), since is stable, For case (iii), follows from Lemma 3.3 of [19] (note that, proved in [18], if has real rank zero, then ). Therefore
| (e 6.253) |
@@
| (e 6.254) |
In general, we have the following statement, which follows immediately from Theorem 4.6. Following the same spirit as the proof of Corollary 6.2, one could alternatively give a direct proof without invoking Theorem 4.6. However, since this approach is already demonstrated in Corollary 6.2, we avoid introducing additional -theoretic technicalities – particularly to spare readers less familiar with -theory.
@@
We now present a counterexample demonstrating that the -dimensional analogue of Theorem 1.3 (i) is false when Note that, in the next proposition, the triple has only a single point in the essential -spectrum. Hence the -theory obstruction vanishes in This example also underscores the significance of Theorems 1.1 and 1.2.
@@
where infimum is taken from all possible commuting triples of selfadjoint operators and all
@@
Choose Identify with a -subalgebra of such that becomes a projection of rank In particular, as a finite rank () projection. Put Define by
| (e 6.263) |
Choose (with )
| (e 6.264) |
@@
| (e 6.277) |
This contradicts (e 6.259). Therefore there is no homomorphisms such that (e 6.269) holds. In other words, there is such that
| (e 6.278) |
where the infimum is taken among all homomorphisms Combining this with (e 6.266), the proposition follows. ∎
@@
This remark was added to the original version of the paper after we had the opportunity to study Chapter 14 of David Mumford’s insightful new book [34], kindly provided by the author. Some terminologies and motivation are drawn from this work.
@@
In quantum theory, macroscopic observables may be modeled as self-adjoint operators on an infinite dimensional separable Hilbert space Let be the unital -subalgebra generated by and (assuming for convenience). Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonal to) a finite dimensional subspace vanish, meaning an external observer would detect nothing. We exculde this case by assuming none of ’s are compact (see also Remark 6.5).
@@
Given a state defined by a unit vector the expected value of observable is (see p.178 of [34]). Mumford introduces Approximately Macroscopically Unique (AMU) states as “near eigenvectors”: for a small
@@
where (see Chapter 14 of [34]). Mumford asked whether AMU exist. As noted in [34], for AMU to be non-empty, it is natural to require the commutators to be small. If the question posed in the introduction (see (e 1.1) and (e 1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one, ensuring the existence of AMU states. However, as observed by Mumford, small commutators alone are insufficient in general. (Proposition 6.4 further shows the complexity of the problem).
@@
In an infinite dimensional separable Hilbert space a self-adjoint operator such as may lack eigenvalues but admit approximate eigenvalues. However, if one allows non-normal states of or that of (where is the ideal of compact operators), then does have eigenstates and eigenvalues in the atomic representation of or (see, for example, Proposition 4.3.10 of [35]). In particular, one should consider those states of which vanish on
For a point in the unit cube of to be a joint approximate eigenvalue, we may test it using the function defined in 2.7. In order to have (for small ), one only needs one vector state (with a unit vector ) such that
| (e 6.279) |
@@
One may expect this can be repeatedly measured. Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (e 6.279) also holds for (in place of ), and this persists. Then we obtain a sequence of () such that (e 6.279) holds for each (in place of ). Note that unit vector states given by has limit points which give states of vanishing on and
@@
Let be the quotient map and be a state of the corona algebra (or that of ). Then the composition provides a quantum state which may be better suitable to this setting. Indeed, one may expect at least for one such quantum state satisfying
| (e 6.280) |
@@
Such measurements, given by states of the form may be called outside measurements. Via the Gelfand-Naimark construction, this state corresponds to a vector state in the atomic representation space yielding
| (e 6.281) |
@@
If inequality (e 6.280) holds whenever (e 6.279) does, then the second condition in (e 1.3) always satisfies (in fact, in this case ).
Let us temporerly propose the following terminology:
A unit vector satisfying (e 6.279) is a synthetic-AMU state, with a synthetic-joint-expected value.
A state of the form satisfying (e 6.280) is an essential synthetic-AMU state, with an essential synthetic-joint-expected value.
@@
Note that any AMU state (with sufficiently small ) is an ()-synthetic-AMU state. Moreover, by Proposition 2.11, the set of synthetic-AMU states is non-empty, when is small (independent of ). The second condition in Theorem 1.1 then may be reformulated as:
@@
(C): all synthetic-joint-expected values lie in a sufficiently small neighborhood of the set of essential synthetic-joint-expected values.
@@
This seems to suggest a physical constraint on macroscopic systems: local measurement should not deviate significantly from outside measurements (or a long-term measurements along an orthonormal basis of the Hilbert space). Under this constraint, when commutators tend to zero, by Theorem 1.1, the quantum system becomes approximable by a classical commutative one.